OpenStudy (anonymous):

Solve the following quadratic equation by using the quadratic formula. 5x²+2x-7=0

OpenStudy (anonymous):

1st multiply coefficients if 1st n last terms ==5 *-7=-35 find products of -35 whose sum or difference=+2 -35=7 *-5 where 7-5=2 u can slplit 5x²+2x-7=0 as 5x²+7x-5x-7=0 x(5x+70-1(5x+7)=0 proceed nxt

OpenStudy (anonymous):

next

OpenStudy (anonymous):

(5x+7)(x-1)=0 noe equate each term =0 to get x values

OpenStudy (anonymous):

x= 7/5, 1

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

just small - difference in x=7/5